# The Oracle Complexity of Simplex-based Matrix Games: Linear Separability and Nash Equilibria

### v3

@@ 4.1 Oracle complexity with a one-sided oracle
Suppose $d>2T+1$. Then for any deterministic algorithm for solving Eq. (2), there exists a $(T+1)\times d$ matrix $A$ satisfying
$\max_{l\in[T+1]}\|A_{l}\|=1~~~\text{and}~~~\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{1}{\sqrt{T+1}}~,$
@@
By fixing some $\gamma_{A}>0$ and setting $T=\lceil\gamma_{A}^{-2}\rceil$, we can restate this result as follows: For any $\gamma_{A}$, and for any algorithm whose interaction with $A$ is captured by a one-sided oracle, there exists a matrix $A$ (with unit rows and margin parameter at least $\gamma_{A}$) such that the required number of iterations to find a linear separator is $\Omega(\gamma_{A}^{-2})$, as claimed in the introduction.
We emphasize that the proof of this particular theorem is a rather straightforward adaptation of existing techniques, and its main purpose is to complete the picture regarding the power of different oracles to solve this matrix game. Nevertheless, the proof provided below illustrates the iterative nature of such constructions, an idea which is also used in the proofs of our other results.
@@
The proof is directly inspired by standard oracle complexity lower bounds for convex Lipschitz optimization due to Nemirovski and Yudin (1983), as well as the standard lower bound proof of the Perceptron algorithm (cf. Shalev-Shwartz and Ben-David, 2014, section 9.5, question 3). The main idea is to construct $A$’s row as mutually orthogonal unit vectors, which are also orthogonal to the algorithm’s queries $\mathbf{w}_{t}$. Therefore, the algorithm can recover the rows one at a time, by querying for a particular row. However, if the number of rows is larger than $T$, there will remain rows orthogonal to $\mathbf{w}_{T+1}$. Therefore, $\mathbf{w}_{T+1}$ will not be a linear separator for $A$. On the other hand, since $A$ has $T+1$ orthogonal rows, it can be shown to be linearly separable with margin $1/\sqrt{T+1}$.
@@
        Set the $l_{t}$-th row of $A_{t}$ to be some arbitrary unit vector, orthogonal to $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$ as well the rows of $A_{t}$ indexed by $\mathcal{L}_{t}$.
@@
    • Feed $\mathcal{A}$ with $A_{t}\mathbf{w}_{t}$ and with the $l_{t}$-th row of $A_{t}$ (as a response for its queries $\mathbf{w}_{t},l_{t}$).
@@
  • Set $A=A_{T}$, and set all rows $l\notin\mathcal{L}_{T}$ of $A$ to be some mutually orthogonal unit vectors which are also orthogonal to $\mathbf{w}_{1},\ldots,\mathbf{w}_{T+1}$.
We note that since $2T+1<d$, the dimensionality is sufficiently high to find orthogonal unit vectors as specified. Moreover, it is easy to verify that because of the orthogonality, it holds that $A_{t}\mathbf{w}_{t}=A\mathbf{w}_{t}$ for all $t\in[T]$: Namely, the responses given to the algorithm are consistent with the oracle responses on the matrix $A$. However, after $T$ iterations, $|\mathcal{L}_{T}|\leq T$, yet there are $T+1$ rows. Therefore, at least one row of $A$ will be chosen to be orthogonal to $\mathbf{w}_{1},\ldots,\mathbf{w}_{T+1}$, and in particular to $\mathbf{w}_{T+1}$. Hence, $A\mathbf{w}_{T+1}$ contains a 0 entry, so
@@
On the other hand, the unit vector $\mathbf{w}:=\frac{1}{\sqrt{T+1}}\sum_{l=1}^{T+1}A_{l}$ (where $A_{l}$ is the $l$-th row of $A$, with the rows being mutually orthogonal unit vectors) satisfies $A\mathbf{w}~=~\frac{1}{\sqrt{T+1}}\cdot\mathbf{1}$, and therefore,
$\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{1}{\sqrt{T+1}}~.$
4.2 Oracle complexity with a two-sided oracle
We now turn to an oracle complexity lower bound for the much stronger two-sided oracle $\mathcal{O}_{2}^{A}$:
@@ 4.2 Oracle complexity with a two-sided oracle
The following holds for some large enough universal constant $c>0$: For any $T>c$, suppose $d,n$ are sufficiently large so that $d>cT$ and $n>cT^{2}\log(T)$. Then for any deterministic algorithm for solving Eq. (2), there exists an $n\times d$ matrix $A$ satisfying
$\max_{l\in[n]}\|A_{l}\|\leq 1~~~\text{and}~~~\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{1}{cT^{3/2}\sqrt{\log(n)}}~,$
@@
As before, we can state this lower bounds in terms of a margin parameter $\gamma_{A}$: Given any small enough $\gamma_{A}$, we can choose $T$ such that $1/(cT^{3/2}\sqrt{\log(n)})\geq\gamma_{A}$, and get that for any algorithm, there exists an $n\times d$ matrix $A$ (for large enough $n,d$) with margin parameter at least $\gamma_{A}$ and rows of norm at most $1$, such that the required number of iterations $T$ to find a linear separator is $\Omega(\log^{-1/3}(n)\cdot\gamma_{A}^{-2/3})$.
We now turn to discuss the proof of Theorem 4.2. We begin by noting that the proof technique of Theorem 4.1 is of no use here, as it is based on revealing the rows of $A$ to the algorithm one at a time. This is not possible with a two-sided oracle, which allows us (for instance) to compute an arbitrary weighted combination of all rows of $A$ using a single query. A second difficulty is that we consider a very simple class of homogeneous bilinear functions, which means that any lower bound necessarily has to be of this form (as opposed to more general min-max optimization problems, where for lower bounds we can use functions with a richer structure). To handle these difficulties, we introduce a different proof technique, which still forces the algorithm to discover information about $A$ in an incremental manner, but in terms of certain vector outer products rather than rows. Specifically, we will utilize the following randomized construction of $A$:
Construction 1.
@@
    • Compute algorithm queries $\mathbf{p}_{t},\mathbf{w}_{t}$ (based on oracle outputs received so far).
    • Let $\mathbf{u}_{t}$ be some unit vector in $\mathbb{R}^{d}$, which is orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{u}_{t-1}$ and to $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$.
    • Let $\mathbf{v}_{t}=\beta(I-M_{t}M_{t}^{\top})\boldsymbol{\xi}_{t}$, where (1) the columns of $M_{t}\in\mathbb{R}^{n\times s_{t}}$, $s_{t}\leq 2t$ are an orthogonal basis for $\mathbf{v}_{0},\ldots,\mathbf{v}_{t-1},\mathbf{p}_{1},\ldots,\mathbf{p}_{t}$, and (2) $\boldsymbol{\xi}_{t}$ is an independent standard Gaussian random vector in $\mathbb{R}^{n}$.
@@
Note that by construction, $\mathbf{v}_{t}$ is a standard Gaussian random vector in $\mathbb{R}^{n}$, scaled by $\beta$ and projected on the subspace orthogonal to $\mathbf{v}_{0},\ldots,\mathbf{v}_{t-1},\mathbf{p}_{1},\ldots,\mathbf{p}_{t}$ (hence, $\mathbf{v}_{t}$ is orthogonal to all these vectors). This is possible by the assumption $2T+1\leq n$, which implies that $M_{t}$ is always a “thin” matrix representing the basis of a strict subspace of $\mathbb{R}^{n}$. Similarly, the assumption $2T+1\leq d$ implies that choosing $\mathbf{u}_{t}\in\mathbb{R}^{d}$ as orthogonal to 2t12t-1 given vectors (as described above) is indeed possible. The choice of a Gaussian distribution is crucially used in order to control the sign and magnitudes of various quantities associated with the matrix, as will be seen later in the proof.
For the construction to be useful, we need to ensure that the oracle responses as defined above (using intermediate matrices $A_{t}$) are all consistent with the same fixed matrix $A$ in hindsight. This is formalized in the following lemma, whose proof easily follows from the orthogonality properties in the construction, as detailed in the appendix.
@@
$A\mathbf{w}_{T+1}~=~\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}_{T+1}~.$
A second crucial requirement for the construction is that the resulting matrix $A$ is linearly separable with some margin. This is formalized in the following lemma:
Lemma 4.4 (Separator exists).
For $A$ as defined in Construction 1, it holds that
$\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{\alpha}{\sqrt{T+1}}~.$
The proof of Lemma 4.4 notes that $A\mathbf{w}=\frac{\alpha}{\sqrt{T+1}}\mathbf{1}$ for $\mathbf{w}:=\frac{1}{\sqrt{T+1}}\sum_{t=1}^{T+1}\mathbf{u}_{t}$, which is a unit vector following the orthonormality properties in the construction, and hence the maximum is at least this value, as detailed in the appendix.
A final consistency requirement for the construction is that we need to choose the $\alpha,\beta$ parameters appropriately, to satisfy the constraint that each row of $A$ must have norm at most $1$ (at least with high probability, which implies that a suitable choice of $A$ exists). This can be ensured via the following lemma:
Lemma 4.5 (Rows of $A$ are bounded).
Suppose $A$ is constructed as in Construction 1. Then for any $\delta\in(0,1)$ such that $2\alpha^{2}+8T\log\left(\frac{2Tn}{\delta}\right)\beta^{2}\leq 1$ it holds with probability at least $1-\delta$ that $\max_{l\in[n]}\|A_{l}\|\leq 1$, where $A_{l}$ is the $l$-th row of $A$.
The proof of the lemma appears in the appendix, and follows from Gaussian concentration properties. With these consistency components in place, the main task now is to show that the algorithm’s output $\mathbf{w}_{T+1}$ cannot be a linear separator for $A$ (or at least, with high probability over the randomized choice of $A$, which implies that a suitable $A$ exists). This is formalized in the following proposition:
@@
Suppose $A$ is constructed as in Construction 1, and that $\frac{4\alpha}{\beta}\leq\frac{1}{\sqrt{T}}$. Then for any $\delta\in(0,1)$, if $T\sqrt{\frac{80\log(2T/\delta)}{n}}\leq\frac{1}{4}$, then with probability at least $1-\delta-\exp\left(T\log(2n)-\frac{n}{32}\right)$ over the choice of $\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{T}$, it holds that
@@
The formal proof (which appears in the appendix) is rather involved. At a high level, we use Gaussian concentration properties and an $\epsilon$-net argument, to show that with high probability over the randomized construction, it holds for any $\mathbf{w}$ that the vector $\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}$ contains an entry which is upper bounded by a certain expression, which we then prove to be non-positive. Therefore, $\min_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}$ is non-positive for all $\mathbf{w}$. Combining the proposition with the lemmas above, and choosing the $\alpha,\beta$ parameters appropriately, Theorem 4.2 follows, as detailed in the appendix.
Remark 3 (Number of oracle queries vs. number of matrix-vector multiplications).

### v4

@@ 4 Warm up: Linear separability with a one-sided oracle
Suppose $d>4T+32\log(4T)$. Then for any deterministic algorithm for solving Problem (1) for some $p\in[1,2]$, there exists a $(T+1)\times d$ matrix $A$ satisfying $\|A\|_{p\to\infty}=\max_{l\in[T+1]}\|A_{l}\|_{q}\leq 1$ and
$\max_{\mathbf{w}:\|\mathbf{w}\|_{p}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{1}{\sqrt{T+1}}\cdot\begin{cases}1&\text{if~~}p=2\\ \frac{1}{64\log^{3/2}(4d)}&\text{if~~}p\in[1,2)\end{cases}~,$
@@
We note that in the theorem, we require $\|A\|_{p\to\infty}\leq 1$, to ensure that $|A_{l}\mathbf{w}|\leq 1$ for any row $l\in[n]$ and $\mathbf{w}$ in the unit $\ell_{p}$ ball, which is the relevant normalization of $A$ in this problem (of course, one can also re-scale $\|A\|_{p\to\infty}$ to any other fixed value without loss of generality).
By fixing some $\gamma_{A}>0$ and setting $T=\lfloor 2^{-12}\gamma_{A}^{-2}\log^{-3}(4d)\rfloor,~d=\lceil 4T+32\log(4T)\rceil$, we can restate this result as follows: For any $\gamma_{A}$, and for any algorithm whose interaction with $A$ is captured by a one-sided oracle, there exists a matrix $A$ (with unit rows in $q$-norm and margin parameter at least $\gamma_{A}$) such that the required number of iterations to find a linear separator is $\tilde{\Omega}(\gamma_{A}^{-2})$, and even $\Omega(\gamma_{A}^{-2})$ when $p=2$, as claimed in the introduction. This reassures the optimality of the Perceptron algorithm in this model of computation, and more generally, of its non-Euclidean generalizations (up to log factors).
We emphasize that the proof of this particular theorem is a rather straightforward adaptation of existing techniques, and its main purpose is to complete the picture regarding the power of different oracles to solve this matrix game. Nevertheless, the proof below illustrates the iterative nature of such constructions, concretely via iterative Gaussian projections, an idea that also appears in the proofs of our other results to follow.
Remark 4.2.
It should be possible to remove the log factor $1/\log^{3/2}(d)$ that appears in the stated result for $p\in[1,2)$, at the cost of a more involved analysis. The main purpose of our presentation here is to set the stage for the two-sided results to follow, and to prove a polynomial separation of the rates of one- vs. two-sided oracles.
@@
The proof is directly inspired by standard oracle complexity lower bounds for convex Lipschitz optimization due to Nemirovski and Yudin (1983), as well as the standard mistake lower bound proof of the Perceptron algorithm (cf. Shalev-Shwartz and Ben-David, 2014, Section 9.5, Question 3). The main idea is to construct $A$’s row as mutually orthogonal unit vectors, which are also orthogonal to the algorithm’s queries $\mathbf{w}_{t}$. Therefore, the algorithm can recover the rows one at a time, by querying for a particular row. However, if the number of rows is larger than $T$, there will remain rows orthogonal to $\mathbf{w}_{T+1}$. Therefore, $\mathbf{w}_{T+1}$ will not be a linear separator for $A$. On the other hand, since $A$ has $T+1$ orthogonal rows, it can be shown to be linearly separable with margin roughly $1/\sqrt{T}$.
@@
        Let $(A_{t})_{l_{t}}=\frac{1}{\|\Pi_{t}(\boldsymbol{\xi}_{t})\|_{q}}\Pi_{t}(\boldsymbol{\xi}_{t})$ be the normalized projection of an independent Gaussian $\boldsymbol{\xi}_{t}\sim\mathcal{N}(\mathbf{0},I_{d})$, where $\Pi_{t}:\mathbb{R}^{d}\to\mathrm{span}((\mathbf{w}_{i})_{i=1}^{t},(A_{t})_{\mathcal{L}_{t}})^{\perp}$ is the projection onto the subspace orthogonal to $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$ as well the rows of $A_{t}$ indexed by $\mathcal{L}_{t}$.
@@
    • Feed $\mathcal{A}$ with $A_{t}\mathbf{w}_{t}$ and $(A_{t})_{l_{t}}$ (as a response for its queries $\mathbf{w}_{t},l_{t}$).
@@
  • Set $A=A_{T}$, and set all rows $l\notin\mathcal{L}_{T}$ of $A$ to be some mutually orthogonal vectors which are also orthogonal to $\mathbf{w}_{1},\ldots,\mathbf{w}_{T+1}$ with unit $q$-norm.
We note that the dimensionality $d$ is sufficiently high to ensure that the rows of $A$ are mutually orthogonal unit vectors as specified. Moreover, it is easy to verify that because of the orthogonality, it holds that $A_{t}\mathbf{w}_{t}=A\mathbf{w}_{t}$ for all $t\in[T]$: Namely, the responses given to the algorithm are consistent with the oracle responses on the matrix $A$. However, after $T$ iterations, $|\mathcal{L}_{T}|\leq T$, yet there are $T+1$ rows. Therefore, at least one row of $A$ will be chosen to be orthogonal to $\mathbf{w}_{1},\ldots,\mathbf{w}_{T+1}$, and in particular to $\mathbf{w}_{T+1}$. Hence, $A\mathbf{w}_{T+1}$ contains a 0 entry, so
@@
On the other hand, the vector $\mathbf{w}:=\frac{1}{\|\sum_{t=1}^{T+1}A_{t}\|_{p}}\sum_{t=1}^{T+1}A_{t}$ has unit $\ell_{p}$ norm, and satisfies by the orthogonality property for any $l\in[T+1]:$
$(A\mathbf{w})_{l}=\frac{1}{\|\sum_{t=1}^{T+1}A_{t}\|_{p}}\sum_{t=1}^{T}A_{t}^{\top}A_{l}=\frac{\|A_{l}\|_{2}^{2}}{\|\sum_{t=1}^{T+1}A_{t}\|_{p}}~.$ (2)
Recalling that $A$’s rows are projected Gaussians, and therefore are Gaussians themselves, [^2] To be precise, they are Gaussians with random covariance matrices. we can invoke Gaussian concentration arguments to lower bound the ratio above. This results in the following lemma, whose proof is deferred to Section 7.
Lemma 4.3.
AA constructed as above satisfies with probability at least $\frac{1}{2}$ that for all $l\in[T+1]:$
$\frac{\|A_{l}\|_{2}^{2}}{\|\sum_{t=1}^{T+1}A_{t}\|_{p}}\geq\frac{1}{\sqrt{T+1}}\cdot\begin{cases}1&\text{if~~}p=2\\ \frac{1}{64\log^{3/2}(4d)}&\text{if~~}p\in[1,2)\end{cases}~.$
Overall, by the probabilistic method, there is a suitable matrix $A$ so that Eq. (2) and Lemma 4.3 ensure that
$\max_{\mathbf{w}:\|\mathbf{w}\|_{p}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{1}{\sqrt{T+1}}\cdot\begin{cases}1&\text{if~~}p=2\\ \frac{1}{64\log^{3/2}(4d)}&\text{if~~}p\in[1,2)\end{cases}~.$
[…]
$\max_{\mathbf{w}:\|\mathbf{w}\|_{p}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~,$ (3)
with respect to the much stronger two-sided oracle $\mathcal{O}_{2}^{A}$.
@@
The following holds for some large enough universal constant $C>0$ and any $p\in[1,\infty]$: For any $T>C$, suppose $n,d$ are sufficiently large so that $d>CT$ and $n>CT^{2}\log(T)$. Then for any deterministic algorithm for solving Problem (3), there exists a matrix $A\in\mathbb{R}^{n\times d}$ satisfying $\|A\|_{p\to\infty}\leq 1$ and
$\max_{\mathbf{w}:\|\mathbf{w}\|_{p}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}$~ $\geq~\frac{1}{CT(\sqrt{T\log(nd)}+\log(nd))}\cdot\begin{cases}1&\text{if~~}p\in[1,2]\\ 1/\sqrt{\log(d)}&\text{if~~}p\in(2,\infty]\end{cases}~,$
@@
Remark 5.2.
Note that by setting $n,d$ in the result above so that $\sqrt{T\log(nd)}\gg\log(nd)$, the lower bound scales as $\Omega(1/T^{3/2}\sqrt{\log(nd)})$ for $p\in[1,2]$, and as $\Omega(1/T^{3/2}\log(nd))$ for $p\in(2,\infty]$. In any case, a crude bound that holds regardless of the values of $n,d$ and $p$ is $\Omega(1/T^{3/2}\log^{3/2}(nd))$.
As before, we can state this lower bound in terms of the suboptimalty. By fixing small enough $\epsilon>0$ and choosing $T,n,d$ such that the lower bound in the theorem is at least $\epsilon$, we can restate the bound as follows: If the number of iterations $T$ is less than $\Omega(\log^{-1/3}(nd)\epsilon^{-2/3})$ when $p\in[1,2]$ or less than $\Omega(\log^{-2/3}(nd)\epsilon^{-2/3})$ if $p\in(2,\infty]$, there exists a $n\times d$ matrix $A$ such that the optimal value is at least $\epsilon$, yet the algorithm’s output has value less than 0, hence is suboptimal by at least $\epsilon$.
We now turn to discuss the proof of Theorem 5.1. We begin by noting that the proof technique of Theorem 4.1 is of no use here, as it is based on revealing the rows of $A$ to the algorithm one at a time. This is not possible with a two-sided oracle, which allows the algorithm to compute an arbitrary weighted combination of all rows of $A$ using a single query. A second difficulty is that we consider a very simple class of homogeneous bilinear functions, which means that any lower bound necessarily has to be of this form (as opposed to more general min-max optimization problems, where for lower bounds we can use functions with a richer structure). To handle these difficulties, we introduce a different proof technique, which still forces the algorithm to discover information about $A$ in an incremental manner, but in terms of certain vector outer products rather than rows. Specifically, we will utilize the following randomized construction of $A$:
Lower Bound Construction 1.
@@
    • Compute algorithm $\mathcal{A}$’s queries $\mathbf{p}_{t},\mathbf{w}_{t}$ (based on oracle outputs received so far).
    • Let $\mathbf{v}_{t}=\beta\cdot\Pi_{t}(\boldsymbol{\xi}_{t})$ be the scaled projection of an independent Gaussian $\boldsymbol{\xi}_{t}\sim\mathcal{N}(\mathbf{0},I_{n})$, where $\Pi_{t}:\mathbb{R}^{n}\to\mathrm{span}(\mathbf{v}_{1},\ldots,\mathbf{v}_{t-1},\mathbf{p}_{1},\ldots,\mathbf{p}_{t})^{\perp}$ is the projection onto the subspace orthogonal to $\mathbf{v}_{1},\ldots,\mathbf{v}_{t-1},\mathbf{p}_{1},\ldots,\mathbf{p}_{t}$.
    • Let $\mathbf{u}_{t}=\frac{1}{\|\Pi^{\prime}_{t}(\boldsymbol{\xi}^{\prime}_{t})\|_{2}}\Pi^{\prime}_{t}(\boldsymbol{\xi}^{\prime}_{t})$ be the normalized projection of an independent Gaussian $\boldsymbol{\xi}^{\prime}_{t}\sim\mathcal{N}(\mathbf{0},I_{d})$ (also independent of $\boldsymbol{\xi}_{t}$) where $\Pi^{\prime}_{t}:\mathbb{R}^{d}\to\mathrm{span}(\mathbf{u}_{1},\ldots,\mathbf{u}_{t-1},\mathbf{w}_{1},\ldots,\mathbf{w}_{t})^{\perp}$ is the projection onto the subspace orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{u}_{t-1},\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$.
@@
Note that by construction, $\mathbf{v}_{t}$ is a standard Gaussian random vector in $\mathbb{R}^{n}$, scaled by $\beta$ and projected on the subspace orthogonal to $\mathrm{span}(\mathbf{v}_{0},\ldots,\mathbf{v}_{t-1},\mathbf{p}_{1},\ldots,\mathbf{p}_{t})$, hence, $\mathbf{v}_{t}$ is orthogonal to all these vectors. This is a non-trivial projection (i.e., not zero) due to the assumption $2T+1\leq n$. Similarly, the assumption $2T+1\leq d$ ensures that the construction of $\mathbf{u}_{1},\dots,\mathbf{u}_{T+1}\in\mathbb{R}^{d}$ results in an orthonormal set (with probability $1$). The choice of a Gaussian distribution is crucially used in order to control the sign and magnitudes of various quantities associated with the resulting matrix, as will be seen later in the proof.
For the construction to be valid, we need to ensure that the oracle responses as defined above, using intermediate matrices $A_{t}$, are all consistent with the same fixed matrix $A$ in hindsight. This is formalized in the following lemma, whose proof easily follows from the orthogonality properties in the construction, as detailed in the proof section.
@@
$A\mathbf{w}_{T+1}~=~\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}_{T+1}~.$ (4)
A second crucial requirement for the construction is that the resulting matrix $A$ defines a game with some (sufficiently large) positive value. This is formalized in the following lemma:
Lemma 5.4.
For any $\delta\in(0,1)$ and $A$ as defined in Construction 1, there is an absolute constant $c>0$ such that with probability at least $1-\delta:$
$\max_{\mathbf{w}:\|\mathbf{w}\|_{p}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~c\,\alpha\cdot\begin{cases}d^{\frac{1}{2}-\frac{1}{p}}/\sqrt{T}&\text{if~~}p\in[1,2]\\ d^{\frac{1}{2}-\frac{1}{p}}/\sqrt{T\log(d/\delta)}&\text{if~~}p\in(2,\infty]\end{cases}~.$
The proof of Lemma 5.4 is based on noting that $A\mathbf{w}=\frac{\alpha}{\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\|_{p}}\mathbf{1}$ for $\mathbf{w}:=\frac{1}{\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\|_{p}}\sum_{t=1}^{T+1}\mathbf{u}_{t}$, which is a vector in the $\ell_{p}$ ball, and hence the maximum over the ball is at least $\frac{\alpha}{\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\|_{p}}$. We lower bound this value with high probability, based on Gaussian properties of the construction.
A final consistency requirement for the construction is that we need to choose the $\alpha,\beta$ parameters appropriately, to satisfy the constraint that the operator norm $\|A\|_{p\to\infty}$ (equivalently, the largest $\ell_{q}$-norm over $A$’s rows where $1/p+1/q=1$) is at most $1$, at least with high probability, implying that a suitable choice of $A$ exists. This can be ensured via the following lemma:
Lemma 5.5.
For any $\delta\in(0,1)$ and $A$ as defined in Construction 1, there is an absolute constant $C^{\prime}>0$ such that with probability at least $1-\delta-\exp(-d/48):$
$\|A\|_{p\to\infty}\leq C^{\prime}\cdot d^{\frac{1}{2}-\frac{1}{p}}\sqrt{\log(nd/\delta)}\left[\alpha+\beta\sqrt{T}+\beta\sqrt{\log(nd/\delta)}\right]~.$
Thus, as long as $\alpha,\beta,\delta$ satisfy
$d^{\frac{1}{2}-\frac{1}{p}}\sqrt{\log(nd/\delta)}\left[\alpha+\beta\sqrt{T}+\beta\sqrt{\log(nd/\delta)}\right]\leq\frac{1}{C^{\prime}}~,$
it holds with probability at least $1-\delta-\exp(-d/48)$ that $\|A\|_{p\to\infty}\leq 1$.
The proof of the lemma appears in the proof section, and follows from Gaussian concentration properties as well. With these consistency components in place, the main task now is to show that the algorithm’s output $\mathbf{w}_{T+1}$ cannot output a vector with positive value (at least with high probability over the randomized choice of $A$, which implies that a suitable $A$ exists). This is formalized in the following proposition:
@@
Suppose AA is constructed as in Construction 1, and that $4\alpha/\beta\leq 1/\sqrt{T}$. Then for any $\delta\in(0,1)$, if $T\sqrt{80\log(2T/\delta)/n}\leq\frac{1}{4}$, then with probability at least $1-\delta-\exp\left(T\log(2n)-n/32\right)$ over the choice of $\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{T}$, it holds that
@@
The formal proof of Proposition 5.6 is rather involved. At a high level, we use Gaussian concentration properties and an $\epsilon$-net argument, to show that with high probability over the randomized construction, it holds for any $\mathbf{w}$ that the vector $\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}$ contains an entry which is upper bounded by a certain expression, which we then prove to be non-positive. Therefore, $\min_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}$ is non-positive for all $\mathbf{w}$.
We can now assemble the claims outlined throughout this section, and set the parameters $\alpha,\beta$ appropriately in order to prove Theorem 5.1:
Proof of Theorem 5.1.
We construct the matrix $A$ as in Construction 1. Examining the conditions in Lemmas 5.4, 5.5 and Proposition 5.6, and ignoring log factors momentarily, we see that in order for all of them to be applicable, we must ensure $\alpha\lesssim\beta/\sqrt{T}$ (Proposition 5.6), satisfy $\beta\lesssim d^{1/p-1/2}/\sqrt{T}$ and $\alpha\lesssim d^{1/p-1/2}$ (Lemma 5.5), while the game value guarantee is $\approx\alpha d^{1/2-1/p}/\sqrt{T}$ (Lemma 5.4). Thus, to make the game value as large as possible, under the constraints we should choose $\beta\approx d^{1/p-1/2}/\sqrt{T}$ and $\alpha\approx\beta/\sqrt{T}\approx d^{1/p-1/2}/T$, thus leading to a value gap of roughly $\alpha d^{1/2-1/p}/\sqrt{T}\approx 1/T^{3/2}$.
More formally, pick $\delta=1/8$ in Lemmas 5.4, 5.5 and Proposition 5.6, and for the universal constant $C^{\prime}$ given by Lemma 5.5, we choose $\beta=\frac{1}{4C^{\prime}d^{1/2-1/p}(\sqrt{T\log(8nd)}+\log(8nd))}$, as well as $\alpha=\frac{\beta}{4\sqrt{T}}=\frac{1}{16C^{\prime}d^{1/2-1/p}\sqrt{T}(\sqrt{T\log(8nd)}+\log(8nd))}$. It is easily verified that with this choice, as well as the theorem assumptions, the conditions in Lemmas 5.4, 5.5 and Proposition 5.6 are all satisfied, and with a union bound, the resulting matrix $A$ satisfies simultaneously with some positive probability $\|A\|_{p\to\infty}\leq 1$ and $\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}_{T+1}\leq 0$, as well as
$\max_{\mathbf{w}:\|\mathbf{w}\|_{p}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}$ $\geq c\,\alpha\cdot\begin{cases}d^{\frac{1}{2}-\frac{1}{p}}/\sqrt{T}&\text{if~~}p\in[1,2]\\ d^{\frac{1}{2}-\frac{1}{p}}/\sqrt{T\log(d/\delta)}&\text{if~~}p\in(2,\infty]\end{cases}$
$=\frac{c}{16C^{\prime}T(\sqrt{T\log(8nd)}+\log(8nd))}\cdot\begin{cases}1&\text{if~~}p\in[1,2]\\ 1/\sqrt{\log(8d)}&\text{if~~}p\in(2,\infty]\end{cases}~.$
Hence, by the probabilistic method, a suitable fixed matrix $A$ satisfying all of the above must exist, and the result simplifies to the statement in Theorem 5.1 for sufficiently large $C>0$. ∎
Remark 5.7 (Number of oracle queries vs. number of matrix-vector multiplications).